Divide 12 into 2 parts such that the sum of their squares is 80
Answer by LinnW(1048)

Start with 6^2 + 6^2 = 36 + 36 = 72,
next 5^2 + 7^2 = 25 + 49 = 74
4^2 + 8^2 = 16 + 64 = 80
So the answer is 4 and 8